### v1

# Prior Knowledge Accelerate Variance Computing

Abstract
Variance is a basic metric to evaluate the degree of data dispersion, and it is also frequently used in the realm of statistics. However, due to the computing variance and the large dataset being time-consuming, there is an urge to accelerate this computing process. The paper suggests a new method to reduce the time of this computation, it assumes a scenario in which we already know the variance of the original dataset, and the whole variance of this merge dataset could be expressed in the form of addition between the original variance and a remainder term. When we want to calculate the total variance after this adds up, the method only needs to calculate the remainder to get the result instead of recalculating the total variance again, which we named this type of method as PKA(Prior Knowledge Acceleration). The paper mathematically proves the technique’s effectiveness in PKA of variance, and the conditions for this method to accelerate properly.
keywords:
Variance Computing, Acceleration, Prior Knowledge
1 Introduction and Related Works
Variance not only had been widely used in the realm of data analysis like ANOVA[2] but also applied to some probability models in machine learning, such as Linear Gaussian Model[4], Bayesian Regression or estimated the differences as well as contribution between each model in an ensemble model[3].
Yet variance is important in that realm, the calculating variance of it with big data size would be a problem. Except for using various approaches to approximate real value[6], or using matrix block computation to accelerate the process of its general form covariance directly[1], there still exists another solution-adding prior knowledge.
As for accelerating general variance calculation, the book of Sheldon M. Ross, suggests an efficient way to update the variance without recalculating the whole variance [5] when knowing the sample variance of original data. But it only discusses the case when adding data only holds one sample. If just violently using this method for each increment of samples, the accumulated complexity would far exceed the original one.
2 Methodology
In this paper, we further extend the scenario to a more general situation and try to figure out why and when this acceleration is effective. When knowing some existing variable, the expression could be further simplified and accelerate the whole calculating process by using prior knowledge, the paper defined this realm or method as PKA(Prior Knowledge Acceleration).
2.1 PKA(Prior Knowledge Acceleration)
First, we defined a sequence $D$ representing the whole dataset after the extra data gets added to the original one in eq.(1). $D1$ represent the original dataset, and $D2$ is the added one. And both of them are not empty, otherwise, the problem we discussed would be meaningless.
$\forall D=\{x_{1},\dots,x_{N}\}$
$D1=\{x_{1},\dots,x_{N1}\},D2=\{x_{N1+1},\dots,x_{N}\}$ (1)
2.1.1 PKA for Sample Variance
The eq.(2) shows the main concept of this accelerating method, it assumes the variance of the whole dataset could be as formed in its original variance and an unknown remainder.
$\sigma_{D}^{2}=\sigma_{D1}^{2}+R$ (2)
The paper defined the size of the set $D$ as $N$, and defined size $N1$ and $N2$ for its subsets $D1$, and $D2$.
$N=N1+N2$ (3)
Getting rid of $\sigma_{D1}^{2}$ in both side of eq.(2), it could transform to eq.(4).
$R=\sigma_{D}^{2}-\sigma_{D1}^{2}$ (4)
We could extract the term $\sigma_{1}$ and $\sigma_{2}$ out of the expression from the definition(seen details from eq.(5) to eq.(7).) of the population variance $\sigma^{2}$ of $D$, and then we could get the simplified result in eq.(8).
$\sigma_{D}^{2}=\frac{\sum_{x_{i}\in D}(x_{i}-\mu)^{2}}{N}$ (5)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D1}(x_{j}-\mu)^{2}+\sum_{x_{k}\in D2}(x_{k}-\mu)^{2})$ (6)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D1}(x_{j}-\mu_{1}+\mu_{1}-\mu)^{2}+\sum_{x_{k}\in D2}(x_{k}-\mu_{2}+\mu_{2}-\mu)^{2})$ (7)
$\sigma_{D}^{2}=\frac{1}{N}(N1\sigma_{D1}^{2}+N2\sigma_{D2}^{2}+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})$ (8)
Since we need to get the expression of the Remainder $R$, we could put back the eq.(8) into eq.(2). Hence, the final form of the Remainder could explicate in beneath(seen in eq.(10)):
$R=\frac{1}{N}(N1\sigma_{D1}^{2}+N2\sigma_{D2}^{2}+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})-\sigma_{D1}^{2}$ (9)
$R=\frac{1}{N}(N2(\sigma_{D2}^{2}-\sigma_{D1}^{2})+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})$ (10)
As for PKA in mean value when knowing $\mu_{1}$ and $\mu_{2}$, the expression could seen in eq.(11).
$\mu=\frac{N1\mu 1+N2\mu 2}{N}$ (11)
2.1.2 PKA for Population Variance
What if we are calculating sample variance instead of the population one? We could get back to eq.(7) and substitute the variable in it to get eq.(12).
$S_{D}^{2}=\frac{1}{N-1}(\sum_{x_{j}\in D1}(x_{j}-\bar{X_{j}}+\bar{X_{j}}-\bar{X})^{2}+\sum_{x_{k}\in D2}(x_{k}-\bar{X_{k}}+\bar{X_{k}}-\bar{X})^{2})$ (12)
$=\frac{1}{N-1}((N1-1)S_{D1}^{2}+(N2-1)S_{D2}^{2}+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))$ (13)
And using a similar process in eq.(13), we could finally get the expression of sample variance in eq.(15).
$R=\frac{1}{N-1}((N1-1)S_{D1}^{2}+(N2-1)S_{D2}^{2}$
$+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))-S_{D1}^{2}$ (14)
$R=\frac{1}{N-1}((N2-1)S_{D2}^{2}-N2S_{D1}^{2})$
$+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))$ (15)
2.2 Proof of PKA’s effectiveness in Population Variance
In the paper, we assume the computation time as a unit for one single addition is $u_{a}$, and $u_{m}$ for one single multiplication. $M$ is the number of multiplications needed to compute, and $A$ is the number of additions. Based on the definition of TABLE 1, we could get the computing time $t_{p}$ of $\sigma_{D}^{2}$ in PKA could be written as a form in eq.(16).
Table 1: Computational Times for Parameters
Term A M
 $\sigma_{D}^{2}$ 2N-1 N
$\sigma_{D1}^{2}$ 2N1-1 N1
$\sigma_{D2}^{2}$ 2N2-1 N2
$\mu$ N-1 1
$\mu$ in PKA 1 3
$\mu_{1}$ N1-1 1
$\mu_{2}$ N2-1 1
$R$ 5 4
Total $R$ N+2N2+2 N2+9
$\sigma_{D}^{2}$ in PKA N+2N2+2 N2+10
$t_{p}=u_{a}(N+2N2+2)+u_{m}(N2+10)$ (16)
Next step, we need to compare the relation between the $t_{p}$ and the computing time of calculating variance $t$ directly.
$t-t_{p}=u_{a}((2N-1)-(N+2N2+2))+u_{m}(N-(N2+10))$ (17)
$t-t_{p}=u_{a}(N1-N2-3)+u_{m}(N1-10)$ (18)
When PKA is effectiveness means $t-t_{p}>0$, we further simplified it and defined an accelerating factor $\lambda_{a}$, the factor gets larger means PKA is more effective in calculating variance.
$\lambda_{a}=\frac{t}{t_{p}}=\frac{u_{a}(N1-N2-3)}{u_{m}(N1-10)}<1$ (19)
Considering the extreme conditions, when $N1$ approaches infinity, which means we hold an extremely large original dataset in that case, the limit of factor $\lambda_{a}$ is:
$\lambda_{a}=\lim_{N1\to\infty}\frac{u_{a}(N1-N2-3)}{u_{m}(N1-10)}=\frac{u_{a}}{u_{m}}<1$ (20)
This explains, with enough original data, the PKA method is established and effective in all circumstances. As for increasing N2, the factor is also less than the one illustrated in eq.(20).
$\lambda_{a}=\lim_{N2\to\infty}\frac{u_{a}(N1-N2-3)}{u_{m}(N1-10)}=-\infty<1$ (20)
But how about N1 and N2 all approach infinity? In that case, the limit of factors is also fulfilled about the conditions.
$\lambda_{a}=\lim_{(N1,N2)\to(\infty,\infty)}\frac{u_{a}(N1-N2-3)}{u_{m}(N1-10)}=-\frac{-3u_{a}}{\infty}<1$ (20)
To get the case that PKA is not work, we could reconstruct the term and get out of $N1$ or $N2$, and get the conditions that PKA does nothing or worse than directly calculating definition, N1 and N2 both show cannot be too small compare to its counterpart(N1 compare to N2, N2 compare to N1) based on the running environment.
$N1<\frac{u_{a}N2+3u_{a}+10u_{m}}{u_{a}-u_{m}}$ (21)
$N2<\frac{u_{a}(N1-3)+u_{m}(N1-10)}{u_{a}}$ (22)
3 Experiment
As for validating that the method is truly accelerating and has its value, the experiment is settled in a standard Kaggle environment, making it easy to replicate. We generate 25,$0$ random samples in $D1$ and 25,$0$ random samples in $D2$, and those generated data all follow uniform distribution(other distributions also fined, it would not affect the computational time of variance). Based on the previous description, when $N1=N2$ is in the same situation as N1 and N2 both approximate positive infinity and it’s the case that the PKA has better performance compared to the baseline. Kaggle environment holds $\lambda_{a}=-5.9864e-05$ with $u_{a}=1.2433e-06,u_{m}=1.2463e-06$, which is satisfy the condition.
Figure 1: The performance of PKA
As Fig.1 shows, the PKA performance is the one labeled ”accelerated” in the plot, which shows smaller computational time compared to baseline, and when the sample size gets larger this tendency becomes more apparent. When the total sample size N increases to 50,$0$, the PKA reduces $75.60\%$ computing times.
4 Conclusion
In general, the PKA method for calculating variance the article suggests holds an improvement in the computational burden of the variance calculations with the assistance of the knowledge of the variance of the original dataset. The PKA technique finds quicker computation when additional information is incorporated, making it beneficial in situations, especially when processing large volumes of datasets during data analysis.