### v4

# Prior Knowledge Accelerate Variance Computing

Abstract
Variance is a basic metric to evaluate the degree of data dispersion, and it is also frequently used in the realm of statistics. However, due to the computing variance and the large dataset being time-consuming, there is an urge to accelerate this computing process. The paper suggests a new method to reduce the time of this computation, it assumes a scenario in which we already know the variance of the original dataset, and the whole variance of this merge dataset could be expressed in the form of addition between the original variance and a remainder term. When we want to calculate the total variance after this adds up, the method only needs to calculate the remainder to get the result instead of recalculating the total variance again, which we named this type of method as PKA(Prior Knowledge Acceleration). The paper mathematically proves the effectiveness of PKA in variance calculation, and the conditions for this method to accelerate properly.
keywords:
Variance Computing, Acceleration, Prior Knowledge
1 Introduction and Related Works
Variance not only had been widely used in the realm of data analysis like ANOVA[2], but also frequently applied to some probability models in machine learning, such as Linear Gaussian Model[4], Bayesian Regression or estimated the differences as well as contribution between each model in an ensemble model[3].
Yet variance is important in those realms, the calculation of variance and its variants in large data size would be a problem. Except for using various approaches to approximate real value[6], or using matrix block computation to accelerate the computing process directly[1], there still exists another solution-adding prior knowledge.
As for accelerating general variance calculation, the book of Sheldon M. Ross, suggests an efficient way to update the variance without recalculating the whole variance when knowing the sample variance of original data[5]. But it only discusses the case when adding data only holds one sample. If just violently using this method for each increment of samples, the accumulated complexity would far exceed the original one.
2 Methodology
In this paper, we further extend this scenario to a more general situation and try to figure out why and when this acceleration is effective. When knowing some existing variable, the expression could be further simplified and accelerate the whole calculating process by using prior knowledge, the paper defined this realm of methods we called as PKA(Prior Knowledge Acceleration). In this paper, the research only discussing knowing previous variance value for calculations, not including the mean value or other metrics.
2.1 PKA(Prior Knowledge Acceleration)
First, we defined a sequence $D$ representing the whole dataset after the extra data gets added to the original one in eq.(1). $D1$ represent the original dataset, and $D2$ is the added one. And both of them are not empty, otherwise, the problem we discussed would be meaningless. Those sequences are not sets in the definition, which means we are not just simply defined $D1$ and $D2$ are independent to each other.
$\forall D=\{x_{1},\dots,x_{N}\},D1=\{x_{1},\dots,x_{N1}\},D2=\{x_{N1+1},\dots,x_{N}\}$ (1)
2.1.1 PKA for Population Variance
The eq.(2) shows the main concept of this accelerating method, it assumes the variance of the whole dataset could be as formed in its original variance and an unknown remainder.
$\sigma_{D}^{2}=\sigma_{D1}^{2}+R$ (2)
The paper defined the size of the data $D$ as $N$, and defined size $N1$ and $N2$ for its sub-sequences $D1$, and $D2$.
$N=N1+N2$ (3)
Getting rid of $\sigma_{D1}^{2}$ in both side of eq.(2), it could transform to eq.(4).
$R=\sigma_{D}^{2}-\sigma_{D1}^{2}$ (4)
We could extract the term $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ out of the expression from the definition(seen details from eq.(5) to eq.(7).) of the population variance $\sigma^{2}$ of $D$, and then we could get the simplified result in eq.(8).
$\sigma_{D}^{2}=\frac{\sum_{x_{i}\in D}(x_{i}-\mu)^{2}}{N}$ (5)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D1}(x_{j}-\mu)^{2}+\sum_{x_{k}\in D2}(x_{k}-\mu)^{2})$ (6)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D1}(x_{j}-\mu_{1}+\mu_{1}-\mu)^{2}+\sum_{x_{k}\in D2}(x_{k}-\mu_{2}+\mu_{2}-\mu)^{2})$ (7)
$\sigma_{D}^{2}=\frac{1}{N}(N1\sigma_{D1}^{2}+N2\sigma_{D2}^{2}+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})$ (8)
Since we need to get the expression of the Remainder $R$, we could put back the eq.(8) into eq.(2). Hence, the final form of the Remainder could explicate in beneath(seen in eq.(10)):
$R=\frac{1}{N}(N1\sigma_{D1}^{2}+N2\sigma_{D2}^{2}+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})-\sigma_{D1}^{2}$ (9)
$R=\frac{1}{N}(N2(\sigma_{D2}^{2}-\sigma_{D1}^{2})+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})$ (10)
As for PKA in mean value when knowing $\mu_{1}$ and $\mu_{2}$, the expression could seen in eq.(11).Therefore, we could using the R add with original population variance to get the updated one.
$\mu=\frac{N1\mu 1+N2\mu 2}{N}$ (11)
2.1.2 PKA for Sample Variance
What if we are calculating sample variance instead of the population one? We could get back to eq.(7) and substitute the variable in it to get eq.(12).
$S_{D}^{2}=\frac{1}{N-1}(\sum_{x_{j}\in D1}(x_{j}-\bar{X_{j}}+\bar{X_{j}}-\bar{X})^{2}+\sum_{x_{k}\in D2}(x_{k}-\bar{X_{k}}+\bar{X_{k}}-\bar{X})^{2})$ (12)
$=\frac{1}{N-1}((N1-1)S_{D1}^{2}+(N2-1)S_{D2}^{2}+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))$ (13)
And using a similar process in eq.(13), we could finally get the expression of $R$ in eq.(15).
$R=\frac{1}{N-1}((N1-1)S_{D1}^{2}+(N2-1)S_{D2}^{2}+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))-S_{D1}^{2}$ (14)
$R=\frac{1}{N-1}((N2-1)S_{D2}^{2}-N2S_{D1}^{2})+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))$ (15)
2.1.3 Proof of PKA’s effectiveness in Population Variance
In the paper, we assume the computation time as a unit for one single addition is $u_{a}$, and $u_{m}$ for one single multiplication. $M$ is the number of multiplications needed to compute, and $A$ is the number of additions. Based on the definition of TABLE 1, we get the computing time $t_{p}$ of $\sigma_{D}^{2}$ in PKA, it could be written as a form in eq.(16).
Table 1: Computational Times for Parameters
Term A M
 $\sigma_{D}^{2}$ 2N-1 N
$\sigma_{D1}^{2}$ 2N1-1 N1
$\sigma_{D2}^{2}$ 2N2-1 N2
$\mu$ N-1 1
$\mu$ in PKA 1 3
$\mu_{1}$ N1-1 1
$\mu_{2}$ N2-1 1
$R$ 5 6
$R$ in $S_{D}^{2}$ 7 6
Total $R$ N+2N2+3 N2+11
$\sigma_{D}^{2}$ in PKA N+2N2+3 N2+12
$S_{D}^{2}$ in PKA N+2N2+5 N2+12
$t_{p}=u_{a}(N+2N2+3)+u_{m}(N2+13)$ (16)
Next step, we need to compare the relation between the $t_{p}$ and the computing time of calculating variance $t$ directly.
$t=u_{a}(2N-1)+u_{m}N$ (17)
When PKA is effectiveness means $t-t_{p}>0$, we further simplified it and defined an accelerating factor $\tau_{a}$, the factor gets smaller means PKA is more effective in calculating variance.
$\tau_{a}=\frac{t}{t_{p}}>1$ (18)
Considering the extreme conditions, when $N1$ approaches infinity, which means we hold an extremely large original dataset in that case, the limit of factor $\tau_{a}$ is:
$\tau_{a}=\lim_{N1\to\infty}\frac{t}{t_{p}}=\lim_{N1\to\infty}\frac{u_{a}(2N-1)+u_{m}N}{u_{a}(N+2N2+3)+u_{m}(N2+12)}=\frac{2u_{a}+u_{m}}{u_{a}}>1$ (19)
But as for increasing N2, the factor is less than the one illustrated in eq.(20), which is conditions that PKA does nothing or worse than directly calculating definition. This explains, compared to the size of added data, the PKA method is only established and effective with enough original data.
$\tau_{a}=\lim_{N2\to\infty}\frac{t}{t_{p}}=\lim_{N2\to\infty}\frac{u_{a}(2N-1)+u_{m}N}{u_{a}(N+2N2+3)+u_{m}(N2+12)}=\frac{2u_{a}+u_{m}}{3u_{a}+u_{m}}<1$ (20)
2.2 PKA compare with Sheldon M. Ross’s Method
By transforming the definition of sample variance, the Sheldon M. Ross suggests a approach for calculating the new variance and mean value after adding one new sample into the data, representing in eq.(21) as well as eq.(22) [5].
$\bar{x}_{j+1}=x_{j}-\frac{\bar{x}_{j}}{j+1}$ (21)
$s_{j+1}^{2}=(1-\frac{1}{j})s_{j}^{2}+(j+1)(\bar{x}_{j+1}-\bar{x}_{j})^{2}$ (22)
According to the definition, we could get $t_{r}$,the time of the Sheldon M. Ross’s Method when adding $N2$ samples with time function $T$ in eq.(24):
$t_{r}=T(\bar{x}_{N1})+N2(T(\bar{x}_{j+1})+T(s_{j+1}^{2}))$ (23)
$T(\bar{x}_{N1})=(N1-1)u_{a}+u_{m},T(\bar{x}_{j+1})=2u_{a}+u_{m},T(s_{j+1}^{2})=4u_{a}+3u_{m}$ (24)
$t_{r}=(N+5N2-1)u_{a}+(4N2-1)u_{m}$ (25)
$t_{p}^{\prime}=t_{p}+(T(R^{\prime})-T(R))=t_{p}+2u_{a}=u_{a}(N+2N2+5)+u_{m}(N2+13)$ (26)
Further comparing the $t_{r}$ and $t_{p}^{\prime}$(the running of PKA in sample variance) in eq.(25) and eq.(26), we could seen that $t_{r}$ only hold a smaller computing time of additions when $N2<2$(which means N2 can only be one), but this equation conflicts with the scenario that $N2$ have to be greater or equal to 2 for having variance. As for multiplications, $t_{r}<t_{p}^{\prime}$ established only when $N2<int(\frac{11}{3})=3$. To summarize those two conditions to ensure $t_{r}$ must smaller then $t_{p}^{\prime}$, N2 have to equal to one. In conclusion, for dealing with an extremely small amount of adding data, Sheldon M. Ross’s method may present better then PKA in Variance, otherwise it not be a good choice.
2.3 General Form of PKA
After discussing PKA in computing variance, we could further consider about how could this sort approaches applied in a more general scale. In general form of PKA, it also defined two datasets or vectors $D1$ and $D2$ and holds other same assumptions. In general PKA, the distinction is that here we defined a function $f$, which is the function the task aim to calculate rather then variance, and also establish a function $g$ to represent what the paper called remainder when calculating variance:
$f(D1,D2)=f(D1)+f(D2)+g(D1,D2)$ (27)
Once we know the prior knowledge $D1$, it become a constant values, which we using $C1$ to distinguish them:
$f(C1,D2)=f(C1)+f(D2)+g(C1,D2)$ (28)
When the PKA have smaller computing time, it is simply means $T(f(D1,D2))>T(f(C1,D2))$. Additionally, the time function $T$ is a linear function that is a linear combination of unit time. Hence further conduct it we can get:
$T(f(D1))-T(f(C1))>T(g(C1,D2))-T(g(D1,D2))$ (29)
Due to there has no need to calculate $D1$, the term in eq.(30) must be positive.
$T(f(D1))-T(f(C1))>0$ (30)
So we could divide the right side in both side get the general form PKA factor $\tau_{a}$ in eq.(31). The condition is exactly same as the PKA in variance, accelerate factor need to be larger than 1.
$\tau_{a}=\frac{T(f(D1))-T(f(C1))}{T(g(C1,D2))-T(g(D1,D2))}$ (31)
For having a deeper understanding about this condition, we could construct a new function called $h$,and we treat $D2$ as constant in the function, the definition of it could seen in beneath:
$z(x)=T(f(x)),h(z)=T(g(f^{-1}(T^{-1}(z),D2))$ (32)
In that case, eq.(29) could convert to:
$z(D1)-z(C1)>h(z(C1))-h(z(D1))$ (33)
The left side of the term is larger than the right side, we multiply a positive number $L$ in the left, and taking an absolute value on both sides, the inequality is still established. This is fulfilled with Lipschitz’s condition [7]. In other words, the saddle point of this function exists. When there is a constraint that restricts D1 in a limited dataset, these two conditions allow the extreme values also exist. $h(x)$ is a function to describe the relationship between the original calculating time and the accelerated one, hence there exist points that PKA works in general situations.
$L|z(D1)-z(C1)|>|h(z(C1))-h(z(D1))|$ (34)
3 Experiment of PKA in Population Variance
As for validating that the method is truly accelerating and has its value, the experiment is settled in a standard Kaggle environment, making it easy to replicate. We generate 25,$0$ random samples in $D1$ and 25,$0$ random samples in $D2$, and those generated data all follow uniform distribution(other distributions also fined, it would not affect the computational time of variance).
Based on the previous description, when $N1=N2$ is in the same situation as N1 and N2 both approximate positive infinity and it’s the case that the PKA has better performance compared to the baseline. The paper taking the mean values of 10,$0$ running time of one single operation and knowing that Kaggle environment holds approximately $\tau_{a}=1.208097$ with $u_{a}=2.223825-07,u_{m}=2.338433e-07$, which is satisfy the condition $\tau_{a}>1$.
Figure 1: The performance of PKA
As Fig.1 shows, the PKA performance is the one labeled ”accelerated” in the plot, which shows smaller computational time compared to baseline, and when the sample size gets larger this tendency becomes more apparent with $N1=N2$. When the total sample size N increases to 50,$0$, the PKA reduces $22.04\%$ to $75.60\%$ computing times. This test is settled in the case that $N1=N2$, hence the paper creating another more precise experiment about it.
But we need a more clear view of the relationship of those variables , the paper making $N1$ and $N2$ range within $[10e3,10e5]$. And getting the mean values 30 times for smoothing the figure, the result could check the figure of $t-t_{p}$ in Fig.2. According to the figure of the surface, the PKA starts to reduce the time of calculating variance roughly around $N1=3e4$ until the end of testing values, indicating is effective when $N1$ is larger than an unknown certain value. Also, the surface shows somehow a linear relationship between N1, N2, and the calculating time, indicating if $N1$ is larger than a threshold, the PKA still working, which proves the previous proof.
Figure 2: The time that PKA method reduce
4 Conclusion
In general, the PKA method for calculating variance the article suggests holds an improvement in the computational burden of the variance calculations with the assistance of the knowledge of the variance of the original dataset. The PKA technique finds quicker computation when additional information is incorporated, making it beneficial, especially when processing large volumes of datasets during data analysis.
But the factor in PKA is holding a potential issue when calculating variance, it assumes the unit time $u_{m}$ and $u_{a}$ are constants. The assumption of PKA in variance is approximately correct when processing large-size data, but further improvements and validations to it may still needed in the future.
Besides, the PKA methods simplify the terms of the known variables, this process is similar to making a shear in matrix multiplication, it is similar to elimination. Traditional approaches like the Gauss-Seidel method could deal with elimination, but they cannot handle solving multi-term simplifications. Hence if there is a numerical method to simplify this process, PKA could also be adopted for solving other much more complex questions without involving manual conduction.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.