### v7

# Prior Knowledge Accelerate Variance Computing

Abstract
The paper introduces Prior Knowledge Acceleration (PKA), a method to speed up variance calculations by leveraging prior knowledge of the original dataset’s variance. PKA enables the efficient updating of variance when adding new data, reducing computational costs by avoiding full recalculations. We derive expressions for both population and sample variance using PKA and compare it to Sheldon M. Ross’s method. Stimulated results show that PKA can reduce calculation time by up to 75.6%, especially when the original dataset is large. PKA offers a promising approach to accelerate variance computations in large-scale data analysis, although its effectiveness depends on assumptions of constant computational time.
keywords:
Statistic Theory, Variance Computing, Acceleration, Prior Knowledge
1 Introduction and Related Works
Variance not only had been widely used in the realm of data analysis like ANOVA [5] but also frequently applied to some probability models in machine learning, such as Linear Gaussian Model [7], Bayesian Regression or estimated the differences as well as contribution between each model in an ensemble model [6].
Yet variance is important in those realms, the calculation of variance and its variants in large data size would be a problem. Except for using various approaches to approximate the real value [12], or using matrix block computation to directly accelerate the computing process [1], there still exists another solution to prior knowledge.
As for accelerating general variance calculation, the book of Sheldon M. Ross, suggests an efficient way to update the variance without recalculating the whole variance when knowing the sample variance of original data [11]. But it only discusses the case when adding data only holds one sample. If just violently using this method for each increment of samples, the accumulated complexity would far exceed the original one.
2 Methodology
In this paper, we further extend this scenario to a more general situation and try to figure out why and when this acceleration is effective. When knowing some existing variable, the expression could be further simplified and accelerate the whole calculating process by using prior knowledge, in that way, if the reduction of computing is computing less with the more time that complex computing graph brought, PKA could get work. The paper defined this realm of methods we called PKA(Prior Knowledge Acceleration). In this paper, the research only discusses knowing previous variance values for calculations, not including the mean value or other metrics. The simplified overview of the computing graph for population variance as an example can be seen in Fig.1.
Figure 1: The Computing Graph for PKA
2.1 PKA(Prior Knowledge Acceleration)
First, we defined a sequence $D$ representing the whole dataset after the extra data gets added to the original one in eq.(1). $D1$ represent the original dataset, and $D2$ is the added one. In addition, both of them are not empty, otherwise, the problem we discussed would be meaningless. Those sequences are not sets in the definition, which means we are not just simply defining $D1$ and $D2$ are independent to each other.
$\forall D=\{x_{1},\dots,x_{N}\},D1=\{x_{1},\dots,x_{N1}\},D2=\{x_{N1+1},\dots,x_{N}\}$ (1)
2.1.1 PKA for Population Variance
The eq.(2) shows the main concept of this accelerating method, it assumes the variance of the whole dataset could be as formed in its original variance and an unknown remainder.
$\sigma_{D}^{2}=\sigma_{D1}^{2}+R$ (2)
The paper defined the size of the data $D$ as $N$, and defined size $N1$ and $N2$ for its sub-sequences $D1$, and $D2$.
$N=N1+N2$ (3)
Getting rid of $\sigma_{D1}^{2}$ in both side of eq.(2), it could transform to eq.(4).
$R=\sigma_{D}^{2}-\sigma_{D1}^{2}$ (4)
We could extract the term $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ out of the expression from the definition(seen details from eq.(5) to eq.(7).) of the population variance $\sigma^{2}$ of $D$, and then we could get the simplified result in eq.(8).
$\sigma_{D}^{2}=\frac{\sum_{x_{i}\in D}(x_{i}-\mu)^{2}}{N}$ (5)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D1}(x_{j}-\mu)^{2}+\sum_{x_{k}\in D2}(x_{k}-\mu)^{2})$ (6)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D1}(x_{j}-\mu_{1}+\mu_{1}-\mu)^{2}+\sum_{x_{k}\in D2}(x_{k}-\mu_{2}+\mu_{2}-\mu)^{2})$ (7)
$\sigma_{D}^{2}=\frac{1}{N}(N1\sigma_{D1}^{2}+N2\sigma_{D2}^{2}+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})$ (8)
Since we need to get the expression of the Remainder $R$, we could put back the eq.(8) into eq.(2). Hence, the final form of the Remainder could explicate in beneath(seen in eq.(10)):
$R=\frac{1}{N}(N1\sigma_{D1}^{2}+N2\sigma_{D2}^{2}+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})-\sigma_{D1}^{2}$ (9)
$R=\frac{1}{N}(N2(\sigma_{D2}^{2}-\sigma_{D1}^{2})+N1(\mu_{1}-\mu)^{2}+N2(\mu_{2}-\mu)^{2})$ (10)
As for the mean value of PKA in variance(it had been widely use due to its simplicity, it’s a special case of PKA that do not has the third term) when knowing $\mu_{1}$ and $\mu_{2}$, the expression could seen in eq.(11). Therefore, we could use the R add with the original population variance to get the updated one.
$\mu=\frac{N1\mu 1+N2\mu 2}{N}$ (11)
2.1.2 PKA for Sample Variance
What if we are calculating sample variance instead of the population one? We could get back to eq.(7) and substitute the variable in it to get eq.(12).
$S_{D}^{2}=\frac{1}{N-1}(\sum_{x_{j}\in D1}(x_{j}-\bar{X_{j}}+\bar{X_{j}}-\bar{X})^{2}+\sum_{x_{k}\in D2}(x_{k}-\bar{X_{k}}+\bar{X_{k}}-\bar{X})^{2})$ (12)
$=\frac{1}{N-1}((N1-1)S_{D1}^{2}+(N2-1)S_{D2}^{2}+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))$ (13)
Using a similar process in eq.(13), consequently, we could finally get the expression of $R$ in eq.(15).
$R=\frac{1}{N-1}((N1-1)S_{D1}^{2}+(N2-1)S_{D2}^{2}+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))-S_{D1}^{2}$ (14)
$R=\frac{1}{N-1}((N2-1)S_{D2}^{2}-N2S_{D1}^{2})+N1(\bar{X_{j}}-\bar{X})^{2}+N2(\bar{X_{k}}-\bar{X}))$ (15)
2.1.3 Proof of PKA’s effectiveness in Population Variance
In the paper, we assume the computation time as a unit for one single addition is $u_{a}$, and $u_{m}$ for one single multiplication. $M$ is the number of multiplications needed to compute, and $A$ is the number of additions. Based on the definition of TABLE 1, we get the computing time $t_{p}$ of $\sigma_{D}^{2}$ in PKA, it could be written as a form in eq.(16).
Table 1: Computational Times for Parameters
Term A M
 $\sigma_{D}^{2}$ 2N-1 N
$\sigma_{D1}^{2}$ 2N1-1 N1
$\sigma_{D2}^{2}$ 2N2-1 N2
$\mu$ N-1 1
$\mu$ in PKA 1 3
$\mu_{1}$ N1-1 1
$\mu_{2}$ N2-1 1
$R$ 5 6
$R$ in $S_{D}^{2}$ 7 6
Total $R$ N+2N2+3 N2+11
$\sigma_{D}^{2}$ in PKA N+2N2+3 N2+12
$S_{D}^{2}$ in PKA N+2N2+5 N2+12
$t_{p}=u_{a}(N+2N2+3)+u_{m}(N2+13)$ (16)
Next step, we need to compare the relation between the $t_{p}$ and the computing time of calculating variance $t$ directly.
$t=u_{a}(2N-1)+u_{m}N$ (17)
When PKA is effectiveness means $t-t_{p}>0$, we further simplified it and defined an accelerating factor $\tau_{a}$, the factor gets smaller means PKA is more effective in calculating variance.
$\tau_{a}=\frac{t}{t_{p}}>1$ (18)
Considering the extreme conditions, when $N1$ approaches infinity, which means we hold an extremely large original dataset in that case, the limit of factor $\tau_{a}$ is:
$\tau_{a}=\lim_{N1\to\infty}\frac{t}{t_{p}}=\lim_{N1\to\infty}\frac{u_{a}(2N-1)+u_{m}N}{u_{a}(N+2N2+3)+u_{m}(N2+12)}=\frac{2u_{a}+u_{m}}{u_{a}}>1$ (19)
But as for increasing N2, the factor is less than the one illustrated in eq.(20), which is a condition that PKA does nothing or worse than directly calculating the definition. This explains that, compared to the size of added data, the PKA method is only established and effective with enough original data.
$\tau_{a}=\lim_{N2\to\infty}\frac{t}{t_{p}}=\lim_{N2\to\infty}\frac{u_{a}(2N-1)+u_{m}N}{u_{a}(N+2N2+3)+u_{m}(N2+12)}=\frac{2u_{a}+u_{m}}{3u_{a}+u_{m}}<1$ (20)
2.2 PKA compare with Sheldon M. Ross’s Method
By transforming the definition of sample variance, Sheldon M. Ross suggests an approach for calculating the new variance and mean value after adding one new sample into the data, representing in eq.(21) as well as eq.(22) [11].
$\bar{x}_{j+1}=x_{j}-\frac{\bar{x}_{j}}{j+1}$ (21)
$s_{j+1}^{2}=(1-\frac{1}{j})s_{j}^{2}+(j+1)(\bar{x}_{j+1}-\bar{x}_{j})^{2}$ (22)
According to the definition, we could get $t_{r}$, the time of the Sheldon M. Ross’s Method when adding $N2$ samples with time function $T$ in eq.(24):
$t_{r}=T(\bar{x}_{N1})+N2(T(\bar{x}_{j+1})+T(s_{j+1}^{2}))$ (23)
$T(\bar{x}_{N1})=(N1-1)u_{a}+u_{m},T(\bar{x}_{j+1})=2u_{a}+u_{m},T(s_{j+1}^{2})=4u_{a}+3u_{m}$ (24)
$t_{r}=(N+5N2-1)u_{a}+(4N2-1)u_{m}$ (25)
$t_{p}^{\prime}=t_{p}+(T(R^{\prime})-T(R))=t_{p}+2u_{a}=u_{a}(N+2N2+5)+u_{m}(N2+13)$ (26)
Further comparing the $t_{r}$ and $t_{p}^{\prime}$(the running of PKA in sample variance) in eq.(25) and eq.(26), we could see that $t_{r}$ only hold a smaller computing time of additions when $N2<2$(which means N2 can only be one), but this equation conflicts with the scenario that $N2$ have to be greater or equal to 2 for having variance. As for multiplications, $t_{r}<t_{p}^{\prime}$ established only when $N2<int(\frac{11}{3})=3$. To summarize those two conditions to ensure $t_{r}$ must be smaller than $t_{p}^{\prime}$, N2 has to equal to one. In conclusion, for dealing with an extremely small amount of adding data, Sheldon M. Ross’s method may present better than PKA in Variance, otherwise it not be a good choice.
2.3 General Form of PKA
After discussing PKA in computing variance, we could further consider how this sort of approach is applied on a more general scale. In the general form of PKA, it defines two datasets or vectors $D1$ and $D2$, and holds other same assumptions. In general PKA, the distinction is that here we defined a function $f$, which is the function the task aims to calculate rather than variance, and also establish a function $g$ to represent what the paper called remainder when calculating variance:
$f(D1,D2)=Af(D1)+Bf(D2)+g(D1,D2)$ (27)
Once we know the prior knowledge $D1$, it become constant values, which we using $C1$ to distinguish them:
$f(C1,D2)=Af(C1)+Bf(D2)+g(C1,D2)$ (28)
When the PKA has a smaller computing time, it simply means $T(f(D1,D2))>T(f(C1,D2))$. Additionally, the time function $T$ is a linear function that is a linear combination of unit time. The $Bf(D_{2})$ term get offset, hence further conduct we can get:
$A\times(T(f(D1))-T(f(C1)))>T(g(C1,D2))-T(g(D1,D2))$ (29)
Due to there being no need to calculate $C1$, the term in eq.(30) must be positive.
$T(f(D1))-T(f(C1))>0$ (30)
So we could divide the right side in both sides to get the general form PKA factor $\tau_{a}$ in eq.(31). The condition is exactly the same as the PKA in variance, accelerate factor needs to be larger than 1.
$\tau_{a}=A\frac{T(f(D1))-T(f(C1))}{T(g(C1,D2))-T(g(D1,D2))}$ (31)
To have a deeper understanding of this condition, we could construct a new function called $h$, and we treat $D2$ as constant in the function, the definition of it could seen in beneath:
$z(x)=T(f(x)),h(z)=T(g(f^{-1}(T^{-1}(z),D2))$ (32)
In that case, eq.(29) could convert to:
$z(D1)-z(C1)>h(z(C1))-h(z(D1))$ (33)
The left side of the inequality is larger than the right, so multiplying both sides by a positive constant $L$, and in here is L=1, and taking the absolute value preserves the inequality in eq.(34). This satisfies Lipschitz’s condition [13], ensuring the existence of a saddle point. When a constraint limits $D1$ in a finite dataset, these conditions guarantee the existence of extreme values. $h(x)$ describes the relationship between the original and accelerated calculation times, showing that PKA works in general situations.
However, when the function values of the entire dataset involve recursive relationships, as seen in the Master Method [3], the PKA method becomes inapplicable. Similarly, ill-functions like oscillations or discontinuities may make the PKA method ineffective, or the function highly out of the linearity(like transcendental functions), or yet the condition about $\tau$ gets satisfied but the size of data does not reach the point which certainly holds extreme values. But in most cases, PKA still could guarantee its acceleration.
$L|z(D1)-z(C1)|>|h(z(C1))-h(z(D1))|$ (34)
3 Experiment of PKA in Population Variance
As for validating that the method is truly accelerating and has its value, the experiment is settled in a standard Kaggle environment, making it easy to replicate. We generate 25,$0$ random samples in $D1$ and 25,$0$ random samples in $D2$, and those generated data all follow uniform distribution(other distributions are fined, it would not effect the computational time of variance).
Based on the previous description, when $N1=N2$ is in the same situation as N1 and N2 both approximate positive infinity and it’s the case that the PKA has better performance compared to the baseline. The paper taking the mean values of 10,$0$ running time of one single operation and knowing that the Kaggle environment holds approximately $\tau_{a}=1.208097$ with $u_{a}=2.223825-07,u_{m}=2.338433e-07$, which satisfies the condition $\tau_{a}>1$.
Figure 2: The performance of PKA
As Fig.2 shows, the PKA performance is the one labeled ”accelerated” in the plot, which shows smaller computational time compared to baseline, and when the sample size gets larger this tendency becomes more apparent with $N1=N2$. When the total sample size N increases to 50,$0$, the PKA reduces $22.04\%$ to $75.60\%$ computing times. This test is settled in the case that $N1=N2$, hence the paper creating another more precise experiment about it.
But we need a more clear view of the relationship of those variables, the paper making $N1$ and $N2$ range within $[10e3,10e5]$. And getting the mean values 30 times for smoothing the figure, the result could check the figure of $t-t_{p}$ in Fig.3. According to the figure of the surface, the PKA starts to reduce the time of calculating variance roughly around $N1=3e4$ until the end of testing values, indicating is effective when $N1$ is larger than an unknown certain value.
Also, the surface somehow shows a linear relationship between N1, N2, and the calculating time, indicating if $N1$ is larger than a threshold, the PKA still working, which fits the previous proof. The tendency of the test is reliable since the test is from 1 to 50,0001. Each size represents a test, which means it is equal to $500001!$ times validations, which is far more than 30 times, no need for p-value involved or further validations.
Figure 3: The time that PKA method reduces
4 Conclusion
In general, the PKA method for calculating variance the article suggests holds an improvement in the computational burden of the variance calculations with the assistance of the knowledge of the variance of the original dataset.
When the factor $\tau_{a}$ fulfills the conditions, the PKA technique finds quicker computation when additional information is incorporated, making it beneficial, especially when processing large volumes of datasets during data analysis.
But the factor in PKA is holding a potential issue when calculating variance, it assumes the unit time $u_{m}$ and $u_{a}$ are constants. The assumption of PKA in variance is approximately correct when processing large-size data, further improvements and validations to it may still needed in the future. Such as analysis of PKA’s effectiveness in a dynamic environment, or trying to use it to calculate other metrics like skewness.
During the experiment, there is another thing that needs to be noticed, due to the Truncation Error [9], an error caused by limits of storage digits in the computer, there is no surprise that PKA holds a larger Truncation Error since it evolves more step to calculate as Fig.1 show, is around $3010e-$ 30 for type float32 which is slightly higher than directly computing the variance(Its Truncation Error is around $3110e-$ 31).
As for potential applications, incorporating PKA into online machine learning models could be a promising direction [14]. These models typically require constant training and parameter updates (such as gradients and loss values) with newly incoming data, and researchers have tried various methods like Expectation–maximization algorithm(EM) [4] or other numerical methods [10] to estimate the variance [15], hence PKA might be a promising way if get further extensions to covariance or directly adopt it when encountering similar tasks that require frequent parameter updates.
The delay in online learning is always an issue that would usher low precision to the models, extra cost for operations, or add up the error within the calculation [8] and even influence timely decision making [2]. PKA could be one solution for it, as an instance, PKA could be seamlessly incorporated into the Mean Squared Error (MSE) calculation due to its similarity to variance. By adopting PKA in online machine learning, the risk of miscalculations caused by delays could be mitigated, leading to more accurate and efficient updates in real-time training. It is a promising way since the result of PKA is an analytical solution albeit the potential limits.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.