### v9

# Extending Sheldon M. Ross’s Method for Efficient Large-Scale Variance Computation

Abstract
The paper introduces Prior Knowledge Acceleration (PKA), a method to speed up variance calculations by leveraging prior knowledge of the variance in the original dataset. PKA enables the efficient updating of variance when adding new data, reducing computational costs by avoiding full recalculation. We derive expressions for both population and sample variance using PKA and compare them to Sheldon M. Ross’s method. Unlike Sheldon M. Ross’s method, the PKA method is designed for processing large data streams online like online machine learning. Simulated results show that PKA can reduce calculation time in most conditions, especially when the original dataset or added one is relatively large. While this method shows promise in accelerating variance computations, its effectiveness is contingent on the assumption of constant computational time.
keywords:
Statistic Theory, Variance Computing, Acceleration, Prior Knowledge
1 Introduction and Related Works
Variance plays a crucial role in data analysis, such as in ANOVA [5], and is widely applied in probability models within machine learning, including the Linear Gaussian Model [7] and Bayesian Regression. It is also used to estimate differences and contributions between models in ensemble learning [6].
Despite its importance, computing variance and its variations for large datasets can be computationally expensive. Except for using various approaches to approximate the real value [14], or using matrix block computation to directly accelerate the computing process [1], there still exists another solution to prior knowledge. However, beyond these computational strategies, leveraging prior knowledge presents another promising direction for optimization.
In the context of online machine learning, where models require continuous training and frequent parameter updates (e.g., gradients and loss values), variance estimation remains a challenge [16]. These online machine learning models typically require constant training and parameter updates (such as gradients and loss values) with newly incoming data, and researchers have tried various methods like Expectation–maximization algorithm(EM) [4] or other numerical methods [12] to estimate the variance [18]. Even though the ELM, refer as the Extreme Learning Machine(a type of model that could avoid frequent updates of parameters), or other alternative methods get proposed, the main issue in traditional online learning still remains [17]. Hence PKA might be a way if get further extensions to covariance or directly adopt it when encountering similar tasks that require frequent parameter updates.
The delay in online learning is always an issue that would usher low precision to the models, incur extra cost for operations, or add up the error within the calculation [9] and even influence timely decision making [2]. PKA could be one solution for it, as an instance, PKA could be seamlessly incorporated into the Mean Squared Error (MSE) calculation due to its similarity to variance. By adopting PKA in online machine learning, the risk of miscalculations caused by delays could be mitigated, leading to more accurate and efficient updates in real-time training. It is a feasible way since the result of PKA is an analytical solution albeit the potential limits.
For accelerating general variance computation to solve those issues, the similar idea occurred in the book of Sheldon M. Ross, suggesting an efficient way to update the variance without recalculating the whole variance when knowing the sample variance of original data [13]. However, his approach only applies to cases where a single new sample is added. If just naively using this method for each increment of samples, the accumulated complexity would far exceed that of the original variance calculation, that is where the edge of PKA stand for, to accelerate when processing batch of added dataset.
2 Methodology
In this paper, we further extend this scenario to a more general situation and try to figure out why and when this acceleration is effective. When knowing some existing variable, the expression could be further simplified and accelerate the whole calculating process by using prior knowledge, in that way, if the reduction of computing is computing less with the more time that complex computing graph(a data structure represent and decide the priority of computations) brought, PKA could get work. The paper defined this realm of methods we called PKA(Prior Knowledge Acceleration). In this paper, the research only discusses knowing previous variance values for calculations, not including the mean value or other metrics. The simplified overview of the computing graph for population variance as an example can be seen in Fig.1.
Figure 1: The Computing Graph for PKA
2.1 PKA(Prior Knowledge Acceleration)
First, we defined a sequence $D$ representing the whole dataset after the extra data gets added to the original one in eq.(1). $D_{1}$ represent the original dataset, and $D_{2}$ is the added one. In addition, both of them are non-empty sets. Those sequences are not sets in the definition, which means we are not just simply defining $D_{1}$ and $D_{2}$ are independent to each other.
$D=(x_{1},x_{2},\dots,x_{N}),\quad D_{1}=(x_{1},x_{2},\dots,x_{N_{1}}),\quad D_{2}=(x_{N_{1}+1},\dots,x_{N})$ (1)
2.1.1 PKA for Population Variance
The eq.(2) shows the main concept of this accelerating method, it assumes the variance of the whole dataset could be as formed in its original variance and an unknown remainder.We express the updated variance as the sum of the original variance and a ”remainder” term, which accounts for the influence of the new data and the shift in the overall mean.
$\sigma_{D}^{2}=\sigma_{D_{1}}^{2}+R$ (2)
The paper defined the size of the data $D$ as $N$, and defined size $N_{1}$ and $N_{2}$ for its sub-sequences $D_{1}$, and $D_{2}$.
$N=N_{1}+N_{2}$ (3)
Getting rid of $\sigma_{D_{1}}^{2}$ in both side of eq.(2), it could transform to eq.(4).
$R\triangleq\sigma_{D}^{2}-\sigma_{D_{1}}^{2}$ (4)
We could extract the term $\sigma_{D_{1}}^{2}$ and $\sigma_{D_{2}}^{2}$ out of the expression from the definition(seen details from eq.(5) to eq.(7).) of the population variance $\sigma^{2}$ of $D$, and then we could get the simplified result in eq.(8).
$\sigma_{D}^{2}=\frac{\sum_{x_{i}\in D}(x_{i}-\mu)^{2}}{N}$ (5)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D_{1}}(x_{j}-\mu)^{2}+\sum_{x_{k}\in D_{2}}(x_{k}-\mu)^{2})$ (6)
$\sigma_{D}^{2}=\frac{1}{N}(\sum_{x_{j}\in D_{1}}(x_{j}-\mu_{1}+\mu_{1}-\mu)^{2}+\sum_{x_{k}\in D_{2}}(x_{k}-\mu_{2}+\mu_{2}-\mu)^{2})$ (7)
$\sigma_{D}^{2}=\frac{1}{N}(N_{1}\sigma_{D_{1}}^{2}+N_{2}\sigma_{D_{2}}^{2}+N_{1}(\mu_{1}-\mu)^{2}+N_{2}(\mu_{2}-\mu)^{2})$ (8)
Since we need to get the expression of the Remainder $R$, we could put back the eq.(8) into eq.(2). Hence, the final form of the Remainder could explicate in beneath(seen in eq.(10)):
$R=\frac{1}{N}(N_{1}\sigma_{D_{1}}^{2}+N_{2}\sigma_{D_{2}}^{2}+N_{1}(\mu_{1}-\mu)^{2}+N_{2}(\mu_{2}-\mu)^{2})-\sigma_{D_{1}}^{2}$ (9)
$R=\frac{1}{N}(N_{2}(\sigma_{D_{2}}^{2}-\sigma_{D_{1}}^{2})+N_{1}(\mu_{1}-\mu)^{2}+N_{2}(\mu_{2}-\mu)^{2})$ (10)
As for the mean value of PKA in variance(it had been widely use due to its simplicity, it’s a special case of PKA that do not has the third term) when knowing $\mu_{1}$ and $\mu_{2}$, the expression could seen in eq.(11). Therefore, we could use the R add with the original population variance to get the updated one.
$\mu=\frac{N_{1}\mu 1+N_{2}\mu 2}{N}$ (11)
2.1.2 PKA for Sample Variance
What if we are calculating sample variance instead of the population one? The method still follows the same structure as in the population variance case, but with slight adjustments to it.We could get back to eq.(7) and substitute the variable in it to get eq.(12).
$S_{D}^{2}=\frac{1}{N-1}(\sum_{x_{j}\in D_{1}}(x_{j}-\bar{X_{j}}+\bar{X_{j}}-\bar{X})^{2}+\sum_{x_{k}\in D_{2}}(x_{k}-\bar{X_{k}}+\bar{X_{k}}-\bar{X})^{2})$ (12)
$\frac{1}{N-1}((N_{1}-1)S_{D_{1}}^{2}+(N_{2}-1)S_{D_{2}}^{2}+N_{1}(\bar{X_{j}}-\bar{X})^{2}+N_{2}(\bar{X_{k}}-\bar{X}))$ (13)
Using a similar process of eq.(9) in eq.(13), consequently, we could finally get the expression of $R$ in eq.(15). We do not claim this formula is completely new in its algebraic effectiveness (see (15)). Rather, our contribution is mainly re-contextualization to the framework and expansion of it. While Chan et al.[3] focus on numerical error and algorithmic stability, our work centers on the epistemic limits of statistical inference when data has been compressed.
$R=\frac{1}{N-1}((N_{1}-1)S_{D_{1}}^{2}+(N_{2}-1)S_{D_{2}}^{2}+N_{1}(\bar{X_{j}}-\bar{X})^{2}+N_{2}(\bar{X_{k}}-\bar{X}))-S_{D_{1}}^{2}$ (14)
$R=\frac{1}{N-1}((N_{2}-1)S_{D_{2}}^{2}-N_{2}S_{D_{1}}^{2})+N_{1}(\bar{X_{j}}-\bar{X})^{2}+N_{2}(\bar{X_{k}}-\bar{X}))$ (15)
2.1.3 Proof of PKA’s effectiveness in Population Variance
In the paper, we assume the computation time as a unit for one single addition is $u_{a}$, and $u_{m}$ for one single multiplication. $M$ is the number of multiplications needed to compute, and $A$ is the number of additions. Based on the definition of Table 1, we get the computing time $t_{p}$ of $\sigma_{D}^{2}$ in PKA, it could be written as a form in eq.(16).
Table 1: Computational Times for Variance
Term A M
 $\sigma_{D}^{2}$ 2N-1 N
[…]
$\mu$ N1N-1 1
$\mu$ in PKA 1 3
$\mu_{1}$ $N_{1}-1$ 1
$\mu_{2}$ $N_{2}-1$ 1
$R$ 5 6
$R$ in $S_{D}^{2}$ 7 6
Total $R$ $N+2N_{2}+3$ $N_{2}+11$
$\sigma_{D}^{2}$ in PKA $N+2N_{2}+3$ $N_{2}+12$
$S_{D}^{2}$ in PKA $N+2N_{2}+5$ $N_{2}+12$
$t_{p}=u_{a}(N+2N_{2}+3)+u_{m}(N_{2}+13)$ (16)
Next step, we need to compare the relation between the $t_{p}$ and the computing time of calculating variance $t$ directly.
$t=u_{a}(2N-1)+u_{m}N$ (17)
When PKA is effectiveness means $t-t_{p}>0$, we further simplified it and defined an accelerating factor $\tau_{a}$, the factor gets smaller means PKA is more effective in calculating variance.
$\tau_{a}\triangleq\frac{t}{t_{p}}>1$ (18)
Considering the extreme conditions, when $N_{1}$ approaches infinity, which means we hold an extremely large original dataset in that case, the limit of factor $\tau_{a}$ is:
$\tau_{a}=\lim_{N_{1}\to\infty}\frac{t}{t_{p}}=\lim_{N_{1}\to\infty}\frac{u_{a}(2N-1)+u_{m}N}{u_{a}(N+2N_{2}+3)+u_{m}(N_{2}+12)}=\frac{2u_{a}+u_{m}}{u_{a}}>1$ (19)
But as for increasing $N_{2}$, the factor is less than the one illustrated in eq.(20), which is a condition that PKA does nothing or worse than directly calculating the definition. This explains that, compared to the size of added data, the PKA method is only established and effective with enough original data.
$\tau_{a}=\lim_{N_{2}\to\infty}\frac{t}{t_{p}}=\lim_{N_{2}\to\infty}\frac{u_{a}(2N-1)+u_{m}N}{u_{a}(N+2N_{2}+3)+u_{m}(N_{2}+12)}=\frac{2u_{a}+u_{m}}{3u_{a}+u_{m}}<1$ (20)
2.2 PKA compare with Sheldon M. Ross’s Method
By transforming the definition of sample variance, Sheldon M. Ross suggests an approach for calculating the new variance and mean value after adding one new sample into the data, representing in eq.(21) as well as eq.(22) [13].(This section using original signs in the book for fast references)
$\bar{x}_{j+1}=x_{j}-\frac{\bar{x}_{j}}{j+1}$ (21)
$s_{j+1}^{2}=(1-\frac{1}{j})s_{j}^{2}+(j+1)(\bar{x}_{j+1}-\bar{x}_{j})^{2}$ (22)
According to the definition, we could get $t_{r}$, the time of the Sheldon M. Ross’s Method when adding $N_{2}$ samples with time function $T$ in eq.(24):
$t_{r}=T(\bar{x}_{N_{1}})+N_{2}(T(\bar{x}_{j+1})+T(s_{j+1}^{2}))$ (23)
$T(\bar{x}_{N_{1}})=(N_{1}-1)u_{a}+u_{m},T(\bar{x}_{j+1})=2u_{a}+u_{m},T(s_{j+1}^{2})=4u_{a}+3u_{m}$ (24)
$t_{r}=(N+5N_{2}-1)u_{a}+(4N_{2}-1)u_{m}$ (25)
$t_{p}^{\prime}=t_{p}+(T(R^{\prime})-T(R))=t_{p}+2u_{a}=u_{a}(N+2N_{2}+5)+u_{m}(N_{2}+13)$ (26)
Further comparing the $t_{r}$ and $t_{p}^{\prime}$(the running of PKA in sample variance) in eq.(25) and eq.(26), we could see that $t_{r}$ only hold a smaller computing time of additions when $N_{2}<2$(which means $N_{2}$ can only be one), but this equation conflicts with the scenario that $N_{2}$ have to be greater or equal to 2 for having variance. As for multiplications, $t_{r}<t_{p}^{\prime}$ established only when $N_{2}<int(\frac{11}{3})=3$. To summarize those two conditions to ensure $t_{r}$ must be smaller than $t_{p}^{\prime}$, $N_{2}$ has to equal to one. In conclusion, for dealing with an extremely small amount of adding data, directly Sheldon M. Ross’s method may present better than PKA in Variance, otherwise it not be a good choice.
2.3 General Form of PKA
After discussing PKA in computing variance, we could further consider how this sort of approach is applied on a more general scale. In the general form of PKA, it defines two datasets or vectors $D_{1}$ and $D_{2}$, and holds other same assumptions. In general PKA, the distinction is that here we defined a function $f$, which is the function the task aims to calculate rather than variance, and also establish a function $g$ to represent what the paper called remainder when calculating variance:
$\forall D_{1},D_{2},f(D_{1},D_{2})\triangleq Af(D_{1})+Bf(D_{2})+g(D_{1},D_{2})$ (27)
Once we know the prior knowledge $D_{1}$, it become constant values, which we using $C1$ to distinguish them:
$\because f(C1,D_{2})\triangleq Af(C1)+Bf(D_{2})+g(C1,D_{2})$ (28)
When the PKA has a smaller computing time, it simply means $T(f(D_{1},D_{2}))>T(f(C1,D_{2}))$. Additionally, the time function $T$ is a linear function that is a linear combination of unit time. The $Bf(D_{2})$ term get offset, hence further conduct we can get:
$\therefore A\times(T(f(D_{1}))-T(f(C1)))>T(g(C1,D_{2}))-T(g(D_{1},D_{2}))$ (29)
Due to there being no demand to calculate $C1$, the term in eq.(30) must be positive.
$\because T(f(D_{1}))-T(f(C1))>0$ (30)
So we could divide the right side in both sides to get the general form PKA factor $\tau_{a}$ in eq.(31). The condition is exactly the same as the PKA in variance, accelerate factor needs to be larger than 1.
$\tau_{a}\triangleq A\frac{T(f(D_{1}))-T(f(C1))}{T(g(C1,D_{2}))-T(g(D_{1},D_{2}))}$ (31)
To have a deeper understanding of this condition, we could construct two new function called $z$ and $h$, and we treat $D_{2}$ as constant in the $h$ function, the definition of it could seen in beneath:
$z(x)\triangleq(T\circ f)(x),h(z)\triangleq(g\circ f^{-1})(T^{-1}(z),D_{2})$ (32)
In that case, eq.(29) could convert to:
$\therefore z(D_{1})-z(C1)>(h\circ z)(C1)-(h\circ z)(D_{1})$ (33)
The left side of the inequality is larger than the right, so multiplying both sides by a positive constant $L$, and in here is L=1, and taking the absolute value preserves the inequality in eq.(34). This satisfies Lipschitz’s condition [15], ensuring the existence of a saddle point. When a constraint limits $D_{1}$ in a finite dataset, these conditions guarantee the existence of extreme values. $h(x)$ describes the relationship between the original and accelerated calculation times, showing that PKA works in general situations.
$\therefore L|z(D_{1})-z(C1)|>|(h\circ z)(C1)-(h\circ z)(D_{1})|$ (34)
However, when the function values of the entire dataset involve recursive relationships, as seen in the Master Method, the PKA method becomes inapplicable. Similarly, ill-functions like oscillations or discontinuities may make the PKA method ineffective, or the function highly out of the linearity(like transcendental functions), or yet the condition about $\tau$ gets satisfied but the size of data does not reach the range where certainly holds extreme values. But in most cases, PKA still could guarantee its acceleration.
2.4 Examples of using PKA
Excepting the variance or mean value(in eq.(11), also a special case of using PKA methods but with $g(D_{1},D_{2}))=0$) we discussed, there are other functions that could decompose in that way for faster computation including covariance, or other resemble instances like Within-Cluster Sum of Squares(WCSS) in K-Means Clustering[11], Sum of Square in ANOVA, etc, also could adopt it. This section of the paper would further illustrate a deeper understanding of how to utilize it instead of merely variance, using PKA in covariance as an example.
2.4.1 PKA for Covariance
Covariance is the more general form of regular variance, it is a good example to show PKA’s utility. According to the definition of covariance, the 2 dimensional covariance of the whole dataset $\text{Cov}_{D}$ could be written as in eq.(35):
$\text{Cov}_{D}=\frac{1}{N}\sum_{i\in D}(x_{i}-\mu_{x,D})(y_{i}-\mu_{y,D})$ (35)
[…]
$\sum_{i\in D_{1}}(x_{i}-\mu_{x,D})(y_{i}-\mu_{y,D})+\sum_{i\in D_{2}}(x_{i}-\mu_{x,D})(y_{i}-\mu_{y,D})$ (36)
Now, we turn to focus on the first term about $D_{1}$, we add a new term and minus it, which is mathematically equivalent, hence we could get the new forms of $(x_{i}-\mu_{x,D})$ as well as $(y_{i}-\mu_{y,D})$ in eq.(37)(38):
$x_{i}-\mu_{x,D}=(x_{i}-\mu_{x,1})+(\mu_{x,1}-\mu_{x,D})$ (37)
$y_{i}-\mu_{y,D}=(y_{i}-\mu_{y,1})+(\mu_{y,1}-\mu_{y,D})$ (38)
Substituting eq.(37)(38) back into eq.(36), it transform into:
$\sum_{i\in D_{1}}\big{(}(x_{i}-\mu_{x,1})+(\mu_{x,1}-\mu_{x,D})\big{)}\cdot\big{(}(y_{i}-\mu_{y,1})+(\mu_{y,1}-\mu_{y,D})$ (39.1)
Expanding eq.(39.1) gives four parts:
$\sum_{i\in D_{1}}(x_{i}-\mu_{x,1})(y_{i}-\mu_{y,1})$ (39.2)
$\sum_{i\in D_{1}}(x_{i}-\mu_{x,1})(\mu_{y,1}-\mu_{y,D})$ (39.3)
$\sum_{i\in D_{1}}(\mu_{x,1}-\mu_{x,D})(y_{i}-\mu_{y,1})$ (39.4)
$\sum_{i\in D_{1}}(\mu_{x,1}-\mu_{x,D})(\mu_{y,1}-\mu_{y,D})$ (39.5)
Now, in eq.(39.2), we could see the first term is exactly the definition of $N_{1}$ times of $Cov_{1}$(covariance of dataset1). As for eq.(39.3), since term $(\mu_{y,1}-\mu_{y,D})$ is a constant, it could get out of sum, and left $\sum_{i\in D_{1}}(x_{i}-\mu_{x,1})$, which is zero based on the definition of mean value, and that is the same for eq.(39.5).
In summary, the form of the first sum about $D_{1}$ is:
$\sum_{i\in D_{1}}(x_{i}-\mu_{x,D})(y_{i}-\mu_{y,D})=N_{1}\cdot\text{Cov}_{1}+N_{1}(\mu_{x,1}-\mu_{x,D})(\mu_{y,1}-\mu_{y,D})$ (40)
Vice versa, get the form of the sum term for $D_{2}$, and merge it with the first one:
$N\cdot\text{Cov}_{D}=\big{(}N_{1}\cdot\text{Cov}_{1}+N_{1}(\mu_{x,1}-\mu_{x,D})(\mu_{y,1}-\mu_{y,D})\big{)}$ (41.1)
$\big{(}N_{2}\cdot\text{Cov}_{2}+N_{2}(\mu_{x,2}-\mu_{x,D})(\mu_{y,2}-\mu_{y,D})\big{)}$ (41.2)
Organizing the expression, and writing it into general form eq.(27), we could finally get the corresponding value of each term:
$f(D)=\text{Cov}_{D},f(D1)=\text{Cov}_{1},f(D2)=\text{Cov}_{2}$ (42.1)
$A=\frac{N_{1}}{N},B=\frac{N_{2}}{N}$ (42.2)
$g(D1,D2)=\frac{N_{1}}{N}(\mu_{x,1}-\mu_{x,D})(\mu_{y,1}-\mu_{y,D})+\frac{N_{2}}{N}(\mu_{x,2}-\mu_{x,D})(\mu_{y,2}-\mu_{y,D})$ (42.3)
After writing covariance in the form of PKA, we could further discuss its effectiveness as we did in variance.
2.4.2 Proof of Effectiveness of PKA for Covariance
Have a recap of the computational time in Table 1, we were treating the time of computing all could be decomposed by multiples (also including divisions) and additions (including subtraction), based on the previous formula in eq.(42.1)(42.2)(42.3) the time of using PKA in covariance is:
Table 2: Computational Times for Covariance
Term A M
$\text{Cov}_{D}$ (Baseline) 5N35N-3 $N+3$
Stats for $D_{2}$ $5N_{2}-3$ $N_{2}+3$
Global Mean Update 22 66
Covariance Update Step 77 88
$\text{Cov}_{D}$ (PKA Total) $5N_{2}+6$ $N_{2}+17$
In Table 2 we could write the total computational time of baseline(direct calculate covariance) $t_{cov}$ and PKA $t_{covp}$ into:
$t_{cov}=(5N-3)u_{a}+(N_{2}+17)u_{m}$ (43)
$t_{covp}=(5N_{2}+6)u_{a}+(N+3)u_{m}$ (44)
To sum up,in PKA method, the factor $\tau_{a}$ for covariance is:
$\tau_{a}=\frac{t}{t_{p}}=\frac{t_{cov}}{t_{covp}}=\frac{(5N-3)u_{a}+(N_{2}+17)u_{m}}{(5N_{2}+6)u_{a}+(N+3)u_{m}}$ (45)
Too fulfilled the condition in eq.(18), it have to be:
$\frac{(5N-3)u_{a}+(N_{2}+17)u_{m}}{(5N_{2}+6)u_{a}+(N+3)u_{m}}>1$ (46)
Reorganizing eq.(46), it could write as eq.(47),and finally could yield the condition for effectiveness about $N_{1}$ in eq.(48):
$(5N_{1}+3)u_{a}+(-N_{1}+14)u_{m}>0$ (47)
$N_{1}>-\frac{3u_{a}+14u_{m}}{5u_{a}-u_{m}}$ (48)
When $5u_{a}-u_{m}>=0$, the condition is always standing since $N_{1}>0$. When $5u_{a}-u_{m}<0$, which means $5u_{a}<u_{m}$. Based on that, we could change the $3u_{a}+14u_{m}$ into $15u_{m}$, which is surely larger than the original term. For the denominator, we could minus $5u_{a}$, leading the whole expression become eq.(49):
$N_{1}>15=-\frac{15u_{m}}{-u_{m}}>-\frac{3u_{a}+14u_{m}}{5u_{a}-u_{m}}$ (49)
3 Stimulated Tests of PKA in Population Variance
As for validating that the method is truly accelerating and has its value, the experiment is settled in a standard Kaggle environment, making it easy to replicate. By utilizing Numpy package in the first test, we generate 25,0000 random samples in $D_{1}$ and 25,0000 random samples in $D_{2}$, and those generated data all follow uniform distribution(because the distributions not effect the computational time of variance).
When $N_{1}=N_{2}$ is in the same situation as $N_{1}$ and $N_{2}$ both approximate positive infinity and it’s the case that the PKA has better performance compared to the baseline. The paper taking the mean values of 10,0000 running time of one single operation and knowing that the Kaggle environment holds approximately $\tau_{a}=1.2080$ with $u_{a}=2.2238-07,u_{m}=2.3384e-07$, which satisfies the condition $\tau_{a}>1$.
Figure 2: The performance of PKA
As Fig.2 shows, the PKA performance is the one labeled ”accelerated” in the plot, which shows smaller computational time compared to baseline, and when the sample size gets larger this tendency becomes more apparent with $N_{1}=N_{2}$. When the total sample size N increases to 50,0000, the PKA reduces $22.04\%$ to $75.60\%$ computing times. This test is settled in the case that $N_{1}=N_{2}$, hence the paper creating another more precise experiment about it.
But we need a more clear view of the relationship of those variables, the paper making $N_{1}$ and $N_{2}$ range within $[10e3,10e5]$ for making the second test. And getting the mean values 30 times for smoothing the figure, the result could check the figure of $t-t_{p}$ in Fig.4. According to the figure of the surface, the PKA starts to reduce the time of calculating variance roughly around $N_{1}=3e4$ until the end of testing values, indicating is effective when $N_{1}$ is larger than an unknown certain value.
Also, the surface somehow shows a linear relationship between $N_{1}$, $N_{2}$, and the calculating time, indicating if $N_{1}$ is larger than a threshold, the PKA still working, which fits the previous proof. The tendency of the test is reliable since the test is from 1 to 50,0001. Each size represents a test, which means it is equal to $\sum_{n=1}^{500001}{n}=\frac{500001+1}{2}\times 500001$ times validations, which is far more than 30 times, no mandate for p-value involved or further validations.
Figure 3: The time that PKA method reduces
During the experiment, there is another thing that needs to be noticed, due to the Truncation Error [10], an error caused by limits of storage digits in the computer, there is no surprise that PKA holds a larger Truncation Error since it evolves more step to calculate as Fig.1 show, is around 10e3010e-30 for type float32 which is slightly higher than directly computing the variance(Its Truncation Error is around 10e3110e-31).
4 Real-time Experiment
The results in stimulated tests were still lacking support when in an online environment, it might be more dynamic in the aspect of delays or computation power, hence the paper utilizes PKA to compute the variance of Individual Household Electric Power Consumption dataset [8] from UCI Datasets to enhance the conclusion. The reason for choosing it is the form of the dataset is a stream record by day, as well as the easy accessibility to the data, making it perfect for testing the PKA method. In this experiment, we only calculate the feature Global_active_power. The research is using Python package zmq for achieving that(create two process bind to it), and combining the conditions that $S_{1}\times S_{2}=\{(N_{1_{1}},N_{2_{1}}),...|N_{1_{1}}\in S_{1},N_{2_{1}}\in S_{2}\}$, which $S_{1}$ and $S_{2}$ simply are sets all equal to 20,200,2000,20000,20000{20,200,2000,20000,20000}, for grid searching the results. The figure we get from the search still indicates the effectiveness of PKA, the larger the original dataset, the better the PKA works(the difference between the two plots,the PKA and baseline(calculate directly), gets larger).
Figure 4: The comparison test in UCI power consumption dataset
5 Conclusion
In general, the PKA method for calculating variance the article suggests holds an improvement in the computational burden of the variance calculations with the assistance of the knowledge of the variance of the original dataset, addressing critical challenges in large-scale and streaming data environments. When the factor $\tau_{a}$ fulfills the conditions, the PKA technique finds quicker computation when additional information is incorporated, making it beneficial, especially when processing large volumes of datasets during data analysis. Our theoretical analysis indicates that while PKA is highly effective when incorporating small to moderate-sized updates into a large dataset, its benefits diminish when the added data size approaches that of the original dataset.
Our findings also show those PKA methods can achieve reduction in computation time under general conditions. But the factor in all PKA methods are holding a potential issue when calculating variance, it assumes the unit time $u_{m}$ and $u_{a}$ are constants. The assumption of PKA in variance is approximately correct when processing large-size data, further improvements and validations to it may still needed in the future. Such as analysis of PKA’s effectiveness in a dynamic environment, or trying to use it to calculate other metrics like skewness.
Competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding Information
Not Applicable
Author Contribution
Not Applicable
Data Availability Statement
The data were generated using np.random.randn with a fixed seed (8086) for reproducibility. Upon publication, the experiment will be available at https://www.kaggle.com/code/spike8086/pka-method and https://www.kaggle.com/code/spike8086/pka-real-time-test.
Research Involving Human or Animals
Not Applicable
Informed Consent
Not applicable.